Solve the problem.
The amount of money that Pablo mkes in a given week varies directly as the number of hours her works that week. In a week in which he works 38 hours, he makes $589. How much will he make in a week in which he workd 46 hours?
: Solve the problem.
The amount of money that Pablo mkes in a given week varies directly as the number of hours her works that week. In a week in which he works 38 hours, he makes $589. How much will he make in a week in which he workd 46 hours?

Answer by checkley77(3394)

38/589=46/X
38X=589*46
38X=27,094
X=27,094/38
X=$713 AMOUNT EARNED FOR A 46 HOUR WEEK.